how many molecules are in 6.0 moles of methane (CH4)
1
Expert's answer
2018-05-25T07:07:12-0400
Answer:
N = n*Na, where n - amount of substance, mole, Na - Avogadro constant, mole^-1
N = 6*6.022×10^23 = 36.132×10^23